OpenStudy (anonymous):

The work of a student to solve a set of equations is shown below: Equation 1: m = 6 + n Equation 2: 2m = 8 + 8n Step 1: −2(m) = −2(6 + n) 2m = 8 + 8n Step 2: −2m = −12 − 2n 2m = 8 + 8n Step 3: −12 − 2n = 8 + 8n Step 4: −20 = 10n Step 5: n = −2

OpenStudy (anonymous):

In which step did the student first make an error?

OpenStudy (alexandervonhumboldt2):

3 he makked an error with sing

OpenStudy (anonymous):

ty

OpenStudy (solomonzelman):

Please refer to the code of conduct, I have told you so many times not to give out answers. \(\bf\color{blu}{\href{http:///openstudy.com/code-of-conduct}{Code~of~Conduct}}\)

OpenStudy (anonymous):

.

OpenStudy (anonymous):

Which of the following ordered pairs represents the solution to the system given below? x − 4y = 7 5x + 9y = 6

OpenStudy (anonymous):

x - 4y = 7 5x + 9y = 6

OpenStudy (solomonzelman):

you can solve the system, it's not hard: \(\large\color{blue}{ x - 4y = 7 }\) \(\large\color{blue}{ 5x + 9y = 6 }\) multiplying the first equation times -5, \(\large\color{blue}{ -5x +20y = -35 }\) \(\large\color{blue}{ 5x ~~+ 9y = 6 }\)

OpenStudy (solomonzelman):

add the equations, (the X's will cancel) and tell me what you get for y.